OpenStudy (anonymous):

If mike pick 4pds of strawberries and his sister ate 3/5 of them how many did she eat

OpenStudy (texaschic101):

3/5 of 4 lbs (lbs stands for pounds) " of " means multiply 3/5 * 4/1 = 12/5 = 2 2/5 so his sister pigged out on 2 and 2/5 lbs of strawberries...lol.

OpenStudy (anonymous):

Thanks